Grade 12th passPhysical Chemistry

Two moles of N2 and two moles of H2 are taken in a closed vessel of 5 litres capacity and suitable conditions are provided for the reaction. when the equil. is reached , it is found that a half mole of N2 is used up. The equil. concentrartion of NH3 is

vaishnawi singh
10 Years agoGrade 12th pass

2 Answers

Bhavya
10 Years ago
N2 + 3H2 ----> 2NH3
Initial moles 2 2 0
Equilibrium moles 2-\alpha 2-3\alpha 2\alpha
Since 0.5 mole of N2 has been used up therefore :
\alpha= 0.5 mole
Moles of NH3 = 1
Concentration of NH3 = moles/volume = 1/5 mol/l
Rishi Sharma
6 Years ago
Dear Student,
Please find below the solution to your problem.

N2 + 3H2 ----> 2NH3
Initial moles 2 2 0
Equilibrium moles 2-α 2-3α 2α
Since 0.5 mole of N2 has been used up therefore :
α= 0.5 mole
Moles of NH3 = 1
Concentration of NH3 = moles/volume = 1/5 mol/l

Thanks and Regards